Sol 6 Summary – March 17th

Crew 190 Sol Summary Report
17March2018
Sol 6

Summary Title:
The Phoenix great escape!

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 6, 08:00

Sol Activity Summary:
8:15 Medic inspection: memory and reflex tests
8:40 Breakfast
9:30 EVA: Mario (photos), Maximilien (soil sampling), Ariane (support), Michael (support)
9:30 Others: Scientific work
11:00 EVA ended; Crew meeting before lunch
12:30 Lunch
14:30 Scientific work: muography (Sophie), cartography of MDRS (Bastien), chemistry/biology (Ariane, Martin, Maximilien, Frederic), GreenHab (Mario), Phoenix operations & troubleshooting (Michael)
18:00 Report redaction
19:00 CapCom

Look Ahead Plan:
Sol 7:

Notes: an EVA is planned in the afternoon, involving Maximilien, Fred, Sophie, Bastien and Mario. Remaining of the day will be dedicated to scientific work.

Sol 7 to Sol 12:

Notes: previsional planning for the second week of the rotation. Very likely to be modified.

Anomalies in work:
We finally get the Phoenix fully working (except the GPS). Unfortunately, Phoenix got trapped in a tiny hole in the soil; we organised (Michael & Bastien) a rescue mission. Phoenix is now back home, a bit scared though.

Weather:
Sunny in the morning. Cloudy in the afternoon. A bit cold.

Crew Physical Status:
A bit tired, hard to wake up.

EVA:
See EVA report.

Reports to be file:
Commander report
Journalist report
EVA#8 request for Sol 7
Engineer report
Pictures

Support Requested:
None

Sol Summary – March 16th

Crew 190 Sol Summary Report
16March2018
Sol 5

Summary Title:
Phoenix awaken! But still a bit buzzy.

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 6, 08:00

Sol Activity Summary:
8:00 Gym
8:15 Medic inspection: memory and reflex tests
8:40 Breakfast
9:30 EVA: Bastien (cartography – 3D mapping), Martin (experiment check), Sophie (muography), Fred (photos)
9:30 Others: Scientific work
11:00 EVA ended; Crew meeting before lunch
12:30 Lunch
14:30 Scientific work: muography (Sophie), video extraction & compilation (Bastien), chemistry/biology (Ariane, Martin, Maximilien, Frederic), GreenHab (Mario), Phoenix troubleshooting (Michael)
18:00 Report redaction
19:00 CapCom

Look Ahead Plan:
Sol 6:

Notes: EVA (Max, Mario, Ariane, Mick) in the morning (scientific work for the others). Scientific work in the afternoon.

Anomalies in work:
Today we tried to get the Phoenix remote controlled rover working; after fixing some battery issues, we noticed that the rover was not able to connect to the (dedicated) router wifi. We had to manually change the clock of the rover it self in order to fix that problem: the rover had the correct date, but not the router, which we cannot fix (probably the internal clock battery of the router is dead). With similar date/times, router and rover are now able to talk to each other. The video transmission is working, but the commands are still bugging (can’t move the wheels, and cams movement is working randomly).

Weather:
Sunny in the morning. Cloudy in the afternoon. Nice weather.

Crew Physical Status:
Good.

EVA:
See EVA report.

Reports to be file:
Commander report
Journalist report
EVA#7 request for Sol6
Engineer report
GreenHab report
HSO report
Pictures

Support Requested:
None

Crew 190 Sol Summary 14Mar2018

[title Sol Summary -March 14th]

Crew 190 Sol Summary Report
14March2018
Sol 3

Summary Title:
Exploration X Science X Pancakes

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 4, 08:00

Sol Activity Summary:
8:00 Gym
8:20 Breakfast
9:10 EVA: Sophie, Mario, Frederic, Michael
9:15 Cleaning and social activity preparation: Maximilien, Martin, Ariane, Bastien
10:55 EVA ended; Crew meeting before lunch
13:00 Lunch
14:30 Scientific work: muongraphy (Sophie), video extraction & compilation (Bastien), chemistry/biology (Ariane, Martin, Maximilien, Frederic), GreenHab (Mario), operations research and academic work (Michael)
18:00 Redaction of the reports
19:00 CapCom

Optimized Look Ahead Plan:
08:00 – 08:20 Gym
08:20 – 09:00 Breakfast
09:30 – 12:30 EVA: Ariane, Bastien, Maximilien, Michael
09:30 – 12:30 Scientific work: Martin (pharmacology), Sophie (muography, telescope), Mario & Frederic (green hab)
12:30 – 14:00 Lunch
14:00 – 18:00 Back to scientific work: Ariane & Maximilien (chemistry), Bastien (cartography), Michael (operations research)
18:00 – 19:00 Redaction of the reports
19:00 Capcom

Forecast:
To be updated March 15 2018

Anomalies in work:
None

Weather:
Cloudy and a bit cold in the morning. Showers in the afternoon, preventing Sophie from telescope activity.

Crew Physical Status:
In increasingly good shape (thanks to the Gym!) and enthusiasm (pancakes).

EVA:
See EVA report. Mainly soil sampling and pictures.

Reports to be file:
Commander report
Journalist report
EVA#4 report
EVA#5 request (Sol4)
Engineer report
GreenHab
HSO report
Astronomy report

Support Requested:
None

Sol Summary – March 13th

Crew 190 Sol Summary Report
13March2018
Sol 2

Summary Title:
First Rendez-vous with the Martian regolith

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 3, 08:00

Sol Activity Summary:
8:00 Gym
8:20 Breakfast
9:30 EVA: Maximilien, Martin, Ariane, Bastien
9:30 Science/Experimental work: Sophie, Mario, Frederic, Michael
11:00 EVA ended; all crew members switch to SC-work.
13:30 Lunch
14:30 Scientific work: telescope (Sophie), cartography (Bastien), chemistry/biology (Ariane, Martin, Maximilien), GreenHab (Mario, Frederic), operations research (Michael)
18:00 Report redaction
19:00 CapCom

Look Ahead Plan: (automatically generated by our optimizer)
Sol 3:

Sol 3 to 7

Anomalies in work:
None

Weather:

Cloudy in the morning. Clear blue sky in the afternoon, enabling telescope activity. Warm.

Crew Physical Status:

In good shape and still enthusiastic.

EVA:
See EVA report. Mainly soil sampling.

Reports to be file:
Commander report
Journalist report
EVA request for Sol3
Engineer report
HSO report

Crew biographies

(Mission Plan)

Support Requested:
None

Sol Summary 12th

Crew 190 Sol Sumary Report
11March2018
Sol 0

Summary Title:
First Rendez-vous with the Martian regolith

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 2, 08:00
New materials arrived, enabling Sophie’s experiments on Muon’s flows

Sol Activity Summary:
First EVAs: (see EVA reports)
9am to 10am: EVA1 with Maximilien, Martin, Martin and Sophie. 11am to 12am: EVA2 with Ariane, Bastien, Mario and I (Michael). Lunch
Crew meeting: 2pm to 3pm
First experiments from 3pm to 6pm

Global Look Ahead Plan: (generated by our automated scheduler)

Sol2:
(ScreenCapture1)
Adaptations: Ariane will participate to the EVA; Mario will switch with Martin.

Sol2 to Sol6
(ScreenCapture2)
Notes: The planning as modelled here does not mention any very short daily activities, such as checking out bacteries evolution (Ariane) or controlling plant grow (Mario). Also, some experiments (e.g. Martin) involve operations in the very beginning and then long period of wait, before final manipulations.

Anomalies in work:
None

Weather:
Clean sky, a bit cloudy. Warm.

Crew Physical Status:
In good shape and still enthusiastic.

EVA:
First outing for all the crew.

Reports to be file:
Commander report
Journalist report
EVA request for Sol 2
Engineer report

Support Requested:
None

Sol Summary – March 11th

Crew 190 Sol Sumary Report
11March2018
Sol 0

Summary Title:
Welcomed by 189 – Learn – Drive

Michael Saint-Guillain (XO)

Mission Status: Ready to start simulation on Sol 1, 07:00

Sol Activity Summary:
Refill of the secondary water tank (550gal): 7am to 11am.
Introduction by Shannon (11h30am to 2pm):
devices etc, power, water, suits, EVAs
Lunch
Discovering the surroundings by feet (3pm to 5h30pm)
EVA (see EVA)

Look Ahead Plan:
Simulation starts on Sol 1 (tomorrow) 7am with:
sport, breakfast (7-8am)
EVA: 1h30 x2, all crew involved (8h30 – 11h30)
Experimental setups (2pm – 5pm)

Anomalies in work:
None

Weather:
Clean sky almost no clouds, hot.

Crew Physical Status:
In good shape and enthusiastic.

EVA:
Discovering the surroundings with Shannon & Atila.
Visit of the "special region".

Reports to be file:
Commander report
Journalist report
EVA request for Sol 1
Engineer report

Support Requested:
None

Sol Summary – March 10th

Crew 189/190 Sol Summary Report 10 Mars 2018

Sol 20/0

Summary Title: Welcoming our Belgian neighbours

Author: Louis Mangin / Maximilien Richald /Michael Saint Guillain

Crew Status: All nominal

Sol Activity Summary:
Hab cleaning in the morning, debriefings, trip to Green Mars, meeting
with university Rover team. Then arrival of crew 190 from Belgium and
knowledge transfer.

Look Ahead Plan: Briefings for crew 190, installation, trip back to
Grand Junction for crew 189.

Anomalies in work: –

Weather: Partly cloudy, very warm.

Crew Physical Status: Fine.

EVA: –

Reports to be file:
Operation report

Support Requested: –

Sol Summary – March 6th

Crew 189 Sol Summary Report 06 Mar 2018

Sol 16

Summary Title: TV stars

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
Today was another special day: a journalist duo came from Washington
to shoot a report for the news on a French national TV. They were very
curious and friendly and should have taken great shots of Lith Canyon
and the station. The reporter followed us wearing a spacesuit during
the EVA. Let’s hope the result looks great!

The afternoon was quieter: human factors experiments, astronomy,
experiments, video editing, demonstrations/interviews for the
journalists…

Look Ahead Plan: Emergency procedures training vol. 2 in the morning,
small EVA in the afternoon to deploy Solar Panels, GPS guidance
system, shoot missing footage for the rotation video

Anomalies in work: –

Weather: Sunny.

Crew Physical Status: Fine.

EVA: Lith Canyon + Battery change

Reports to be filed:
Commander
Journalist
Engineering
Greenhab Officer
EVA report
EVA request
Astronomy report

Support Requested: –

Sol Summary – March 5th

Here is the Crew 189 Sol Summary Report 05 March 2018

Sol 15

Summary Title: Beginning of last martian week

Author: Victoria Da-Poian

Mission Status: All nominal

Sol Activity Summary: We began our last week on Mars. Today’s EVA was dedicated to air sensors experiments and to localization experiment. The EVA was not really successful (cf EVA report) but we still have one week left to work on this! In the Hab today, I was pleased to cook a salad (salad, cucumbers, cherry tomatoes!): it was delicious! I took care of the greenhab as Jérémy was outside and as the previous Crew 175 Greenhab officer, I spent some time enjoying the vegetables smells! Then, we spent time working on our experiments (human factors experiment, virtual reality filming…)

Look Ahead Plan: Tomorrow, we will welcome two French television journalists (TF1). They will spend the day with us from our physical training to our afternoon activities.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Fine

EVA: 11th

Reports to be filed:
Commander
Journalist
Engineering
Greenhab Officer
Astronomer
EVA

Support Requested: None

Cheers,

Victoria
Crew189 Commander